# ReVision: A Dataset and Baseline VLM for Privacy-Preserving Task-Oriented Visual Instruction Rewriting

### v2

@@ 6 Results and Discussion
Despite the strong performance of our ReVision variants, certain limitations hinder further accuracy gains. A primary issue is the loss of fine-grained text details caused by downsampling images to $256\times 256$ resolution, which impairs recognition of critical elements such as ingredient lists or nutritional facts on product packaging. Additionally, the dataset’s lack of explicit reference localization limits the model’s ability to align user intent with specific image regions, resulting in object disambiguation and instruction alignment errors. Future work could address these challenges by incorporating bounding box annotations to provide spatial grounding cues and by processing localized image regions rather than entire downsampled images, reducing information loss in text-heavy visual inputs. This approach aligns with Pali-Gemma’s short-resolution increase technique beyer2024paligemma, which improves fine-grained visual understanding. Nonetheless, our findings reaffirm that task-specific instruction tuning and metadata augmentation markedly enhance multimodal rewriting, supporting scalable and efficient on-device deployment.

### v3

@@ 6 Results and Discussion
Despite the strong performance of our ReVision variants, certain limitations hinder further accuracy gains. A primary issue is the loss of fine-grained text details caused by downsampling images to $256\times 256$ resolution, which impairs recognition of critical elements such as ingredient lists or nutritional facts on product packaging. Additionally, the dataset’s lack of explicit reference localization limits the model’s ability to align user intent with specific image regions, resulting in object disambiguation and instruction alignment errors. Future work could address these challenges by incorporating bounding box annotations to provide spatial grounding cues and by processing localized image regions rather than entire downsampled images, reducing information loss in text-heavy visual inputs. This approach aligns with Pali-Gemma’s short-resolution increase technique Beyer et al. (2024), which improves fine-grained visual understanding. Nonetheless, our findings reaffirm that task-specific instruction tuning and metadata augmentation markedly enhance multimodal rewriting, supporting scalable and efficient on-device deployment.